OpenStudy (anonymous):

The sun of two numbers is -10. Negative three times the first number minus the second number equals 2. Find the numbers by setting up and solving a system of equations

OpenStudy (anonymous):

x + y = 10 -3x - y = 2

OpenStudy (anonymous):

solve for x, and y